# Alchemy: Amplifying Theorem-Proving Capability through Symbolic Mutation

### v1

@@ 3.2 Proof Generation and Theorem Verification
Mutated statements can serve as useful lemmas for theorem-proving only if we can construct proofs that pass the verification of the proof assistant. We construct the entire proof using symbolic rules. Although neural provers and other automated theorem proving (ATP) tools (e.g., hammer) can generate more natural and diverse proofs than rule-based methods, they are compute-intensive and do not guarantee the correctness of the generated proofs. The idea of building a proof block is intuitive. Given that we only make a one-step modification to the statement, transforming the original proof state to a mutated proof state, a logical approach is to reverse the mutation and utilize the original proof to complete the remaining proving process. We use have tactic to restore the modified part of a statement (the original assumption or assertion) by introducing a lemma.
  • have The have tactic enables users to introduce new assumption into the current proof state if they can prove it. Given an assumption $h_{1}:P$ and an implication rule $h_{2}:P\implies Q$, a new assumption $h:Q$ can be added by have h: Q := by  apply  $h_{2}$ at $h_{1}$; exact  $h_{1}$. This tactic is usually used to introduce helpful lemmas when proving a theorem.
In addition to its ability to introduce new assumptions into the proof state, have can be used in both tactic-style proof and term-style proof, which enlarges the margin for theorems to which our method can be applied. Apart from this, the additional have instruction transforms the mutated complex proof state into a canonical proof state. To some extent, this transformation is analogous to constructing an auxiliary point in geometry problems, which we assume will be beneficial for theorem proving in the general domain. Subsequently, we combine the original proof with this lemma to build the proof for the new statement. The details of the implementation of proof generation are depicted in the Appendix C.3. We construct the proof block for each mutated theorem. Then we submit the synthesized theorems to the Lean theorem prover for verification and remove the wrong ones. Details of the verification process are provided in Appendix C.4. Finally, we obtain a set of variants $V=\{v_{i}\}_{i=0}^{n}$ defined by the keyword “example” for each candidate theorem.
@@ 3.3 Model Training
Regarding the synthetic data, we have two observations. At the theorem level, the synthetic data comprises numerous theorems, each with statement distinct from existing theorems. At the state-tactic level, the process of constructing proofs introduces additional state-tactic pairs, primarily centered on rw and apply. Based on these insights, we assume that the synthetic data can serve as an augmented corpus for continual pretraining and supervised finetuning. Specifically, we fine-tune LLMs using the proofstep prediction objective proposed by Polu & Sutskever (2020), utilizing state-tactic pairs derived from both seed theorems and synthetic theorems. Given the current proof state, the model is required to predict the next tactic sequence that contributes to the proving of the target theorem. We utilize the prompt template used by Welleck (2023), as shown in Fig.2.
@@ 4 Experiments
We implement the data-synthesis pipeline described in Section 3 for rw and apply, constructing a set of variants for each candidate theorem in Mathlib. We train the LLMs on a mixture of human-written theorems and synthetic ones. To examine the effectiveness of synthetic data, we evaluate the theorem prover on two benchmarks that are widely adopted by the research community: 1) Test split of Mathlib, which shares the same distributional characteristics as the seed theorems; 2) miniF2F, a challenging benchmark focusing on competition-level problems that exhibits a distinct distribution compared to seed data. The experimental results derived from both benchmarks demonstrate the potential efficacy of our approach.
@@ 4.1 Implementation Details
We choose Mathlib4333commit: 2iufd 3c307701fa7e9acbdc0680d7f3b9c9fed9081740 which contains 116,695 theorems as the seed data for data-synthesis. Our synthesis pipeline is built upon Leandojo444version: 1.7.1 (Yang et al., 2023), a Python module that enables tracing a specific Lean repository, extracting the state-tactic pairs and abstract syntax trees (ASTs), and interacting with the Lean environment555lean-toolchain: v4.6.0 rc1 (run_tac API). Finding invocable theorems is the most time-consuming step of our pipeline. For rw, the time overhead amounts to 14 days using 4,096 CPU cores666512 CPU nodes, each node has 8 cores and 56GB RAM. For apply, it takes 7 days at this stage using 2,048 CPU cores with a one-hour timeout for each theorem. The substantial time cost is attributed to the $O(n^{2})$ complexity of our algorithm and the memory-intensive characteristics of Leandojo. We believe this overhead could be greatly reduced through a more meticulous implementation. After retrieving the invocable theorems, we construct new statements and proofs for the target theorems in approximately an hour using 24 CPU cores. We then write back the mutated theorems and compile the enlarged repository through lake build, utilizing 2,048 CPU cores. We retrieve the error messages returned by Lean, which can be parsed to locate the wrong theorems. Finally, we trace the enlarged repository on a 96-core machine for 3 days, obtaining the additional state-tactic pairs by parsing the AST of each file.
Model Training
@@
We select Llama-3-8B (Dubey et al., 2024) and deepseek-coder-base-v1.5- 7B (Guo et al., 2024) as our base models. We conduct continual pretraining with the next-token prediction objective for one epoch. Then we fine-tune the models with the proofstep prediction objective (Polu & Sutskever, 2020) for two epochs. All experiments are conducted on $1008\times H$ 100 GPUS. We employ a linear learning rate scheduler with a 3% warm-up period and a maximum learning rate of 2e-5. We set the global batch size to 256 and the cutoff length to 2,048. All models are trained using Deepspeed ZeRO Stage3 (Rajbhandari et al., 2021) and Flash-Attention 2 (Dao, 2023). We utilize the open-sourced codebase Llama-Factory (Zheng et al., 2024) for all training experiments.
Evaluation
@@
We follow the evaluation setting used in Azerbayev et al. (2023). We use the frequently used best-first-search as our search tragedy and set a 10-minute timeout. The search budget can be represented as $N\times S\times T$, where $N$ denotes the number of attempts, $S$ denotes the number of generated tactics per iteration, and $T$ denotes the maximum number of generations. Following Azerbayev et al. (2023), we set N = 1, S = 32 and T = 100. Our evaluation script is modified from an open-source implementation (Welleck, 2023) which is based on vLLM (Kwon et al., 2023) and Leandojo (Yang et al., 2023). We utilize Leandojo Benchmark (Yang et al., 2023) which contains 2,000 theorems as the test split of Mathlib4 and report the results on both the random split and the novel_premises split. We remove the subsets of theorems for both splits that can not be initialized by Leandojo. There remain 1,929 theorems in random split and 1,659 theorems in novel_premises split. We upgrade the tool-chain version of miniF2F (Zheng et al., 2022) to v4.6.0 rc1.
@@ 4.2 Analysis of Synthetic Data
We separately run the synthesis pipeline for these two tactics. For rw, we choose Mathlib theorems as candidate theorems. Additionally, candidate theorems for apply should have at least one explicit assumption. In practice, the synthesis process is divided into two stages. In the first stage, we find the potential invocable theorems for each candidate theorem by running a specific tactic. In the second stage, we construct the new theorems and verify their correctness using the Lean theorem prover. Table 2 shows the number of theorems of different stages. For both tactics, we increase the number of theorems by an order of magnitude (×$\times 25$ for rw and $\times$44 for apply). The conversion ratios from the potential invocable theorems to the outcomes are primarily determined by the method used to construct the new statements and proofs. We believe that a finer implementation could greatly improve the conversion ratio. Figure 3 shows the dynamics of the distribution of mathematical subjects. The rw tactic increases the percentages of Analysis, Ring Algebra, Number Theory, and so on. The apply tactic mainly contributes to the fields of Analysis and Topology. Further information about synthetic data can be found in the Appendix D.
@@
Our method synthesizes a large collection of new theorems utilizing each tactic. Then we combine them with the theorems in Mathlib as the training data for continual pre-training. Our approach also introduces new state-tactic pairs during the theorem-construction process. We write the variants to corresponding lean files and extract additional state-tactic pairs using Leandojo. The synthesized data are categorized primarily based on the employed tactic, specifically rw and apply. Variants and their corresponding state-tactic pairs that appear in the test split of the Leandojo benchmark are removed. Furthermore, the extracted state-tactic pairs are deduplicated according to the invocable theorem (i.e., premise) used in the tactic instruction. Finally, we obtain about 30k data points for each tactic. We combine them with the training set of Leandojo (Mathlib-train) that composes over 200k data points to form the SFT dataset. A detailed description of the deduplication process and training data are presented in the Appendix D.3.
@@ 4.3.1 Main Results
Table 3: Results on Mathlib. tidy: a tactic in Mathlib that uses heuristics to complete a proof. We select the performance of each model solely fine-tuned using Mathlib-train as the main baseline. Mathlib-train + x: the performance of the model pre-trained and fine-tuned on a mixture of Mathlib-train and additional data about x.
@@
We conduct continual pretraining on the augmented lean corpus. Then we fine-tune the LLMs on the mixture of Mathlib-train and additional state-tactic pairs. The training data are grouped by the tactic employed in the additional state-tactic pairs. We evaluate the effectiveness of our method on the challenging Leandojo benchmark and report results on different mixtures of data. As shown in Table 3, our synthetic data consistently improve the theorem-proving capabilities of LLMs. Compared with solely finetuning on the training split of Mathlib, data augmentation for a single tactic demonstrates a beneficial effect on the theorem-proving ability of LLMs. Moreover, the positive impacts of each tactic can be cumulative. Training on the combination of rw variants and apply variants results in a significant performance improvement in the challenging novel_premises split of Leandojo benchmark, where the model is required to use at least one new premise to prove the target theorem (+4.7%, 78 theorems for Llama3-8b; +4.22%, 70 theorems for deepseek-coder-7b-base-v1.5). Our synthetic data still make a certain improvement on the random split, where the performance of models is over-estimated by allowing it to prove many theorems through memorization. In conclusion, the results of the experiment show that simply mutating the seed theorems and introducing state-tactic pairs of a single tactic can relieve the data scarcity problem and enhance the theorem-proving ability of LLMs.
@@ 4.3.3 Influence of the Quantity of SFT Dataset
We deduplicate the synthesized state-tactic pairs of each tactic by the invocable theorem (i.e., premise). Then we obtain about 30k data points for each tactic. To examine the influence of the quantity of the SFT dataset, we compare the performance of Llama-3-8B, trained on different quantities of additional data points, on novel_premises split of Leandojo benchmark. As shown in Fig 4, the selected quantity (30k) achieves a relatively optimal compromise between the performance and overhead. The experimental results also reveal that enlarging the quantity of state-tactic pairs of a single tactic tends to lead to rapid saturation. We assume that the key to continually improving the theorem-proving ability lies in keeping the diversity of tactics during the process of scaling the synthetic data. More details are presented in Appendix D.3.4.
@@ 4.3.4 Analysis of Out-of-distribution Performance
We evaluate Llama-3-8b using the competition-level theorem proving benchmark miniF2F. As shown in Table 5, our synthesized data still helps to improve the theorem-proving ability of LLMs on the out-of-distribution benchmark. The magnitude of this improvement is comparatively smaller than that observed on the in-distribution benchmark. We attribute this discrepancy to the divergence between synthesized tactics and the preferred tactics to prove competition-level problems. Through manual inspection of the correct proofs generated by various LLMs trained on Mathlib-train, we identify a tendency to favor advanced and automated tactics (e.g., simp, omega, linarith, norm_num, etc.). Additionally, we analyze the distribution of tactics used in proved theorems across different data compositions and make the following observations. 1) Data augmentation on a single tactic will increase the model’s preference for the specific tactic. 2) Adjusting the distribution of different tactics within the dataset is promising to improve the theorem-proving ability of LLMs. The entire analysis process is illustrated in Appendix E.2.
@@ 5 Conclusion
We have presented a general data-synthesis framework for the Lean theorem prover, which amplifies the theorem-proving capability of the LLM through symbolic mutation. Our algorithm increases the number of theorems in Mathlib by an order of magnitude and achieves promising results in improving the theorem-proving ability of the LLM. We discuss the limitations of our method in Appendix B. Synthesizing formal theorems is an inherently challenging problem. Our approach, much like ancient alchemy, involves experimenting with a substantial number of theorems in the hope of uncovering valuable “gold”. We aspire for our algorithm and data to serve as a foundation for further research, advancing theorem synthesis from alchemy to chemistry.
@@ Appendix A Background on Lean
Lean is a functional programming language and interactive theorem prover based on dependent type theory. As one of the most popular formal systems, Lean aids mathematicians in formalizing statements and proofs in a semi-auto style and enables them to verify the correctness of each proof step through rigorous type-checking.
Theorem in Lean
To some extent, theorems in Lean can be seen as a special variant of functions in general-purpose programming languages. A theorem consists of a statement and corresponding proof. In Lean, the keyword “theorem”, “example” or “lemma” is used to define the “function”, sometimes followed by a specific function name. The assumption of a statement can be formatted as implicit or explicit arguments, while the assertion of the statement specifies the return type of the function. The proof of the statement can be viewed as the function body, which constructs a proof term with the type specified by the assertion. There are two main proof styles in Lean: term-style and tactic-style. In term-style proofs, theorems are proven using constructive methods. On the other hand, tactic-style proofs sequentially decompose the proof goal using specific tactics. Although tactic-style proofs are less readable, they tend to have shorter proof lengths. Most machine learning-based theorem-proving systems focus on tactic-style proof. The synthesis method proposed by our paper can be applied to both styles.
Tactic
Lean offers various advanced tactics for theorem proving, which set it apart from other formal systems (e.g., Coq, Isabelle). In handwritten proofs, authors tend to guide the reader on building the proof through instructions such as “apply the previous lemma”, “invoke the principle of mathematical induction”, or “simplify the expression”. Similarly, tactics in Lean are used to describe how to construct a proof term incrementally. They help users decompose the proof goal step by step, allowing users to focus on only one proof goal at a time.
Mathlib
Mathlib [^7] https://github.com/leanprover-community/mathlib4 is a comprehensive mathematical library for Lean, largely maintained by the community, which encompasses a broad spectrum of mathematical subjects (e.g., algebra and analysis) and consists of over 120,000 theorems along with their respective axioms and definitions. This extensive knowledge base serves as the primary corpus for neural theorem provers.

### v2

@@ 3.2 Proof Generation and Theorem Verification
Mutated statements can serve as useful lemmas for theorem-proving only if we can construct proofs that pass the verification of the proof assistant. We construct the entire proof using symbolic rules. Although neural provers or other automated theorem proving tools, such as hammer (Böhme & Nipkow, 2010)), can generate more diverse proofs than rule-based methods, they are compute-intensive and do not guarantee the correctness of the generated proofs. The idea of building a proof block is intuitive. Given that we only make a one-step modification to the statement, transforming the original proof state to a mutated proof state, a logical approach is to reverse the mutation and utilize the original proof to complete the remaining proving process. We use have tactic to restore the modified part of a statement (the original assumption or assertion) by introducing a lemma.
  • have. The have tactic enables users to introduce new assumption into the current proof state if they can prove it. Given an assumption $h_{1}:P$ and an implication rule $h_{2}:P\implies Q$, a new assumption $h:Q$ can be added by have h: Q := by  apply  $h_{2}$ at $h_{1}$; exact  $h_{1}$. This tactic is usually used to introduce helpful lemmas when proving a theorem.
In addition to its ability to introduce new assumptions into the proof state, have can be used in both tactic-style proof and term-style proof, which enlarges the margin for theorems to which our method can be applied. Apart from this, the additional have instruction transforms the mutated complex proof state into a canonical proof state. To some extent, this transformation is analogous to constructing an auxiliary point in geometry problems, which we assume will be beneficial for theorem proving in the general domain. Subsequently, we combine the original proof with this lemma to build the proof for the new statement. The details of the implementation of proof generation are depicted in the Appendix C.3. We construct the proof block for each mutated theorem. Then we submit the synthesized theorems to the Lean theorem prover for verification and remove the wrong ones. Details of the verification process are provided in Appendix C.4. Finally, we obtain $\hat{M}$ variants $V=\{v_{j}\}_{j=0}^{\hat{M}-1}$ defined by the keyword “example” for each candidate theorem.
@@ 3.3 Model Training
Regarding the synthetic data, we have two observations. At the theorem level, the synthetic data comprises numerous theorems, each with statement distinct from existing theorems. At the state-tactic level, the process of constructing proofs introduces additional state-tactic pairs, primarily centered on rw and apply. Based on these insights, we assume that the synthetic data can serve as an augmented corpus for continual pretraining (CPT) and supervised finetuning (SFT). Specifically, we fine-tune LLMs using the proofstep prediction objective proposed by Polu & Sutskever (2020), utilizing state-tactic pairs derived from both seed theorems and synthetic theorems. Given the current proof state, the model is required to predict the next tactic sequence that contributes to the proving of the target theorem. We utilize the prompt template used by Welleck (2023), as shown in Fig.2.
@@ 4 Experiments
We implement the data-synthesis pipeline described in Section 3 for rw and apply, constructing a set of variants for each candidate theorem in Mathlib. We train the LLMs on a mixture of human-written theorems and synthetic ones. To examine the effectiveness of synthetic data, we evaluate the theorem prover on two benchmarks that are widely adopted by the research community: 1) Leandojo Benchmark (Yang et al., 2023), which shares the same distributional characteristics as the seed theorems; 2) miniF2F (Zheng et al., 2022), a challenging benchmark focusing on competition-level problems that exhibits a distinct distribution compared to seed data. The experimental results derived from both benchmarks demonstrate the potential efficacy of our approach.
@@ 4.1 Implementation Details
Data-Synthesis. We choose Mathlib4444commit: 3c307701fa7e9acbdc0680d7f3b9c9fed9081740 which contains around 110k theorems as the seed data for data-synthesis. Our synthesis pipeline is built upon Leandojo555version: 1.7.1 (Yang et al., 2023), a Python module that enables tracing a specific Lean repository, extracting the state-tactic pairs and abstract syntax trees (ASTs), and interacting with the Lean environment666lean-toolchain: v4.6.0 rc1 (run_tac API). Finding invocable theorems is the most time-consuming step of our pipeline. For rw, the time overhead amounts to 14 days using 4,096 CPU cores777512 CPU nodes, each node has 8 cores and 56GB RAM. For apply, it takes 7 days at this stage using 2,048 CPU cores with a one-hour timeout for each theorem. The substantial time cost is attributed to the $O(n^{2})$ complexity of our algorithm and the memory-intensive characteristics of Leandojo. We believe this overhead could be greatly reduced through a more meticulous implementation. After retrieving the invocable theorems, we construct new statements and proofs for the target theorems in approximately an hour using 24 CPU cores. We then write back the mutated theorems and compile the enlarged repository through lake build  [^8] https://github.com/leanprover/lean4/blob/master/src/lake/README.md, utilizing 2,048 CPU cores. We retrieve the error messages returned by Lean, which can be parsed to locate the wrong theorems. Finally, we trace the enlarged repository on a 96-core machine for 3 days, obtaining the additional state-tactic pairs by parsing the AST of each file.
@@
Model Training. We select Llama-3-8B (Dubey et al., 2024) and deepseek-coder-base-v1.5- 7B (Guo et al., 2024) as our base models. We conduct continual pretraining with the next-token prediction objective for one epoch. Then we fine-tune the models with the proofstep prediction objective (Polu & Sutskever, 2020) for two epochs. All experiments are conducted on $1008\times H$ 100 GPUS. We employ a linear learning rate scheduler with a 3% warm-up period and a maximum learning rate of 2e-5. We set the global batch size to 256 and the cutoff length to 2,048. All models are trained using Deepspeed ZeRO Stage3 (Rajbhandari et al., 2021) and Flash-Attention 2 (Dao, 2024). We utilize the open-sourced codebase Llama-Factory (Zheng et al., 2024) for all training experiments.
@@
Evaluation. We follow the evaluation setting used in Azerbayev et al. (2024). We use best-first-search as our search tragedy with a 10-minute timeout. The search budget is represented as $attempt\times sample\times step$. Here $attempt$ denotes the number of attempts, $sample$ denotes the number of generated tactics per iteration, and $step$ denotes the maximum number of steps per attempt. We choose $1\times 32\times 100$ as our search setting. The evaluation script is modified from an open-source implementation (Welleck, 2023) which is based on vLLM (Kwon et al., 2023) and Leandojo (Yang et al., 2023). We utilize Leandojo Benchmark (Yang et al., 2023) which contains 2,000 theorems as the test split of Mathlib4 and report the results on both the random split and the novel_premises split. We remove the subsets of theorems for both splits that can not be initialized by Leandojo. There remain 1,929 theorems in random split and 1,659 theorems in novel_premises split. We upgrade the tool-chain version of miniF2F (Zheng et al., 2022) to v4.6.0 rc1.
@@ 4.2 Analysis of Synthetic Data
We separately run the synthesis pipeline for these two tactics. For rw, we choose Mathlib theorems as candidate theorems. Additionally, candidate theorems for apply should have at least one explicit assumption. In practice, the synthesis process is divided into two stages. In the first stage, we find the potential invocable theorems for each candidate theorem by running a specific tactic. In the second stage, we construct the new theorems and verify their correctness using the Lean theorem prover. Table 2 shows the number of theorems of different stages. For both tactics, we increase the number of theorems by an order of magnitude (×$\times 25$ for rw and $\times$44 for apply). The conversion ratios from the potential invocable theorems to the outcomes are primarily determined by the method used to construct the new statements and proofs. We believe that a finer implementation could greatly improve the conversion ratio. Fig.3 shows the dynamics of the distribution of mathematical subjects. The rw tactic increases the percentages of Analysis, Ring Algebra, Number Theory, and so on. The apply tactic mainly contributes to the fields of Analysis and Topology. Further information about synthetic data can be found in the Appendix D.
@@
Our method synthesizes a large collection of new theorems utilizing each tactic. Then we combine them with the theorems in Mathlib as the training data for continual pre-training. Our approach also introduces new state-tactic pairs during the theorem-construction process. We write the variants to corresponding lean files and extract additional state-tactic pairs using Leandojo. The synthesized data are categorized primarily based on the employed tactic, specifically rw and apply. Variants and their corresponding state-tactic pairs that appear in the test split of the Leandojo benchmark are removed. Furthermore, the extracted state-tactic pairs are deduplicated according to the invocable theorem (i.e., premise) used in the tactic instruction. Finally, we obtain about 30k data points for each tactic. We combine them with the training set of Leandojo (Mathlib-train) that composes over 200k data points to form the supervised fine-tuning dataset. A detailed description of the deduplication process and training data are presented in the Appendix D.3.
@@ 4.3.1 Main Results
Table 3: Results on Mathlib. tidy: a tactic in Mathlib that uses heuristics to complete a proof. The results of tidy and GPT4 were reported in Yang et al. (2023). We select the performance of each model solely fine-tuned using Mathlib-train as the main baseline. Mathlib-train + x: the performance of the model pre-trained and fine-tuned on a mixture of Mathlib-train and additional data about x.
@@
We conduct continual pretraining on the augmented lean corpus. Then we fine-tune the LLMs on the mixture of Mathlib-train and additional state-tactic pairs. The training data are grouped by the tactic employed in the additional state-tactic pairs. We evaluate the effectiveness of our method on the challenging Leandojo benchmark and report results on different mixtures of data. As shown in Table 3, our synthetic data consistently improve the theorem-proving capabilities of LLMs. Compared with solely finetuning on the training split of Mathlib, data augmentation for a single tactic demonstrates a beneficial effect on the theorem-proving ability of LLMs. Moreover, the positive impacts of each tactic can be cumulative. Training on the combination of rw variants and apply variants results in a significant performance improvement in the challenging novel_premises split of Leandojo benchmark, where the model is required to use at least one new premise to prove the target theorem (+4.70%, 78 theorems for Llama3-8b; +4.22%, 70 theorems for deepseek-coder-7b-base-v1.5). Our synthetic data still make a certain improvement on the random split, where the performance of models is over-estimated by allowing it to prove many theorems through memorization. In conclusion, the results of the experiment show that simply mutating the seed theorems and introducing state-tactic pairs of a single tactic can relieve the data scarcity problem and enhance the theorem-proving ability of LLMs.
@@ 4.3.3 Influence of the Quantity of SFT Dataset
We deduplicate the synthesized state-tactic pairs of each tactic by the invocable theorem (i.e., premise). Then we obtain about 30k data points for each tactic. To examine the influence of the quantity of the SFT dataset, we compare the performance of Llama-3-8B, trained on different quantities of additional data points, on novel_premises split of Leandojo benchmark. As shown in Fig.4, the selected quantity (30k) achieves a relatively optimal compromise between the performance and overhead. The experimental results also reveal that enlarging the quantity of state-tactic pairs of a single tactic tends to lead to rapid saturation. We assume that the key to continually improving the theorem-proving ability lies in keeping the diversity of tactics during the process of scaling the synthetic data. More details are presented in Appendix D.3.4.
@@ 4.3.4 Analysis of Out-of-distribution Performance
We evaluate Llama-3-8b using the competition-level theorem proving benchmark miniF2F. As shown in Table 5, our synthesized data still helps to improve the theorem-proving ability of LLMs on the out-of-distribution benchmark. The magnitude of this improvement is comparatively smaller than that observed on the in-distribution benchmark. We attribute this discrepancy to the divergence between synthesized tactics and the preferred tactics to prove competition-level problems. Through manual inspection of the correct proofs generated by various LLMs trained on Mathlib-train, we identify a tendency to favor advanced and automated tactics (e.g., simp, omega, linarith, norm_num, etc.). Additionally, we analyze the distribution of tactics used in proved theorems across different data compositions and make the following observations: 1) Data augmentation on a single tactic will increase the model’s preference for the specific tactic; 2) Adjusting the distribution of different tactics within the dataset is promising to improve the theorem-proving ability of LLMs. The entire analysis process is illustrated in Appendix E.2.
@@ 5 Conclusion
We have presented a general data-synthesis framework for the Lean theorem prover, which amplifies the theorem-proving capability of the LLM through symbolic mutation. Our algorithm increases the number of theorems in Mathlib by an order of magnitude and achieves promising results in improving the theorem-proving ability of the LLM. Synthesizing formal theorems is an inherently challenging problem. Our approach, much like ancient alchemy, involves experimenting with a substantial number of theorems in the hope of uncovering valuable “gold”. We aspire for our algorithm to serve as a foundation for further research, advancing theorem synthesis from alchemy to chemistry.
Acknowledgments
This research was supported by the National Natural Science Foundation of China (No. U23B2060, No.62088102). We sincerely thank the Lean Community for providing help about this work. We also appreciate the anonymous reviewers for their helpful comments.
@@ Appendix A Background on Lean
Lean (de Moura et al., 2015) is a functional programming language and interactive theorem prover based on dependent type theory. As one of the most popular formal systems, Lean aids mathematicians in formalizing statements and proofs in a semi-auto style and enables them to verify the correctness of each proof step through rigorous type-checking.
Theorem in Lean. To some extent, theorems in Lean can be seen as a special variant of functions in general-purpose programming languages. A theorem consists of a statement and corresponding proof. In Lean, the keyword “theorem”, “example” or “lemma” is used to define the “function”, sometimes followed by a specific function name. The assumption of a statement can be formatted as implicit or explicit arguments, while the assertion of the statement specifies the return type of the function. The proof of the statement can be viewed as the function body, which constructs a proof term with the type specified by the assertion. There are two main proof styles in Lean: term-style and tactic-style. In term-style proofs, theorems are proven using constructive methods. On the other hand, tactic-style proofs sequentially decompose the proof goal using specific tactics. Although tactic-style proofs are less readable, they tend to have shorter proof lengths. Most machine learning-based theorem-proving systems focus on tactic-style proof. The synthesis method proposed by our paper can be applied to both styles.
Tactic. Lean offers various advanced tactics for theorem proving, which set it apart from other formal systems (e.g., Coq, Isabelle). In handwritten proofs, authors tend to guide the reader on building the proof through instructions such as “apply the previous lemma”, “invoke the principle of mathematical induction”, or “simplify the expression”. Similarly, tactics in Lean are used to describe how to construct a proof term incrementally. They help users decompose the proof goal step by step, allowing users to focus on only one proof goal at a time.
Mathlib. Mathlib [^9] https://github.com/leanprover-community/mathlib4 is a comprehensive mathematical library for Lean, largely maintained by the community, which encompasses a broad spectrum of mathematical subjects (e.g., algebra and analysis) and consists of over 110,000 theorems along with their respective axioms and definitions. This extensive knowledge base serves as the primary corpus for neural theorem provers.